OpenStudy (anonymous):

Solve by simulating the problem. You have a 5-question multiple-choice test. Each question has four choices. You don’t know any of the answers. What is the experimental probability that you will guess exactly three out of five questions correctly?

OpenStudy (anonymous):

Binomial Problem with n = 5 and p(correct) = 1/4 P(x = 3) = 5C3(1/4)^3*(3/4)^2 = binompdf(5,1/4,3) = 0.0879 What do you think of my work correct?

OpenStudy (cwrw238):

probability of guessing right answer for 1 question = 1/4

OpenStudy (cwrw238):

yup - u r right

OpenStudy (anonymous):

okay thanks ^ ^